# by Leveraging CLIP Latent Space

### 07 Feb 2024 v1

@@ Acknowledgments
This work was supported by NSF RI grants #1750082, #2132724, and CPS grant #2038666. The views and opinions of the authors expressed herein do not necessarily state or reflect those of the funding agencies and employers.
@@ Appendix A Preliminaries for T2I Diffusion Models
As evidenced in numerous contemporary studies regarding T2I models, Stable Diffusion (SD) has emerged as a predominant backbone for text-to-image (T2I) models. SD involves training diffusion models in latent space, reversing a forward diffusion process that introduces noise into the image. A notable feature of SD is its integration of cross-attention, facilitating various conditions like text input. Operating in VQVAE latent space, SD not only achieves exceptional generative performance surpassing that in pixel space but also significantly reduces computational demands.
UnCLIP models (such as DALL-E 2) are very similar to the Stable Diffusion. In contrast, the UnCLIP stack takes the modular approach. UnCLIP first trains the diffusion text-to-image to the image prior ($f_{\theta}$) to estimate the image embeddings ($z_{x}$) from the text embeddings ($z_{y}$). In parallel, a UNet-like diffusion image generator ($h_{\phi}$) is trained to generate images ($x$) conditioned on ground truth vision embeddings ($z_{x}$).
@@
$\mathcal{L}_{prior}=\underset{\begin{subarray}{c}t\sim[0,T],\\ z_{x}^{(t)}\sim q(t,z_{x})\end{subarray}}{\operatorname{\mathbb{E}}}\Big{[}||z_{x}-f_{\theta}(z_{x}^{(t)},t,z_{y})||_{2}^{2}\Big{]}.$ (2)
ECLIPSE presents the contrastive learning strategy (Eq. 1) instead of minimizing Eq. 2. The diffusion image generator is trained by following standard $\epsilon$-prediction formulation. Here, $h_{\phi}$ will estimate the ground truth added Gaussian noise $\epsilon\sim N(0,I)$, given the noise image $X^{(t)}$ for all timesteps $t\sim[0,T]$ and input conditions (such as $z_{x}$, $z_{y}$).
$\mathcal{L}_{decoder}=\underset{\begin{subarray}{c}\epsilon\sim N(0,I)\\ t\sim[0,T],\\ (z_{x},~{}z_{y})\end{subarray}}{\operatorname{\mathbb{E}}}\Big{[}||\epsilon-h_{\phi}(x^{(t)},t,z_{x},z_{y})||_{2}^{2}\Big{]}.$ (3)
@@ Appendix B Image-Text Interleaved Training Details
Figure 9: Examples of image-text interleaved training data. Left column shows the input the prior model and right images shows the ground truth corresponding images. Note: these examples are generated from $\lambda$-ECLIPSE for better interpretability.
@@
Method Multi-Concept Finetuning Free Base Model Image Input
@@
Appendix C Implementation Details
Appendix D Extended P-T2I Baselines Comparison
We further expand our comparative analysis of P-T2I methods encompassing a total of 32 approaches including ours and parallel works. Table 5 summarizes them into four crucial aspects: 1) multi-concept support, 2) fine-tuning free, 3) base model types, and 4) the required number of input images. To summarize, $\lambda$-ECLIPSE is the only methodology built on top of the UnCLIP models while supporting multi-concept personalization, fine-tuning free, and only requires a single reference concept image. We detail the comparison below:
Multi-Concept Generation.
Multi-concept generation enables users to integrate multiple personal subjects to generate an image that follows the text prompts and aligns with all the concept visuals. Of the 32 methods analyzed, 14 offer this capability. Ten methods require separate training for each concept and then an additional fusing step for combining the learned concepts. (Mix-of-Show). While only four methods support fast multi-concept personalization. Only a few can learn auxiliary guided information such as canny edge, depth maps, or open-pose and adapt style variation.
@@ Appendix D Extended P-T2I Baselines Comparison
Many models require test-time fine-tuning. Each varies on which part alteration occurs, as early models tend to modify the whole UNet. In contrast, recent models tune a small portion of the cross-attention layers or introduce additional layers performing as adapters. In our analysis of P-T2I models, 13 out of 32 models employ a fine-tuning free approach which enables fast personalization.
@@
A majority of the reviewed models utilize diffusion models, with Stable Diffusion being the predominant choice, spanning versions 1.4, 1.5, 2.1, and XL. Few adapt Imagen (SuTi, Taming) and interestingly Mix-of-show employs ChillOutMix as their pre-trained model, known for its adeptness at preserving realistic concepts like human faces. A unique outlier in this landscape is our $\lambda$-ECLIPSE, the only one that eschews the use of any diffusion prior model.
Easiness to Use (Input and Output).
@@
Ease of use, particularly regarding model input and output, significantly impacts user convenience in subject-driven generalization. A more user-friendly model typically requires a single input image, as opposed to multiple images of the same subject. In our study, 18 models offer P-T2I capabilities with just one input image. In contrast, others often require 4 to 5 images of the subject. Additionally, some models necessitate storage space for learned concepts, ranging from a few hundred kilobytes (e.g., Perfusion, HyperDreamBooth) to several megabytes (e.g., Zip-LoRA). Our model stands out by eliminating the need for individual concept pre-learning or storing any artifacts for P-T2I utilization, offering a streamlined, efficient user experience.
Appendix E More Qualitative Results
In this section, we present an array of extended examples of P-T2I generation, which presents complex compositions that are hard to achieve in $\lambda$-ECLIPSE as well as other models. Figure 10 showcases the complexity level that notably escalates as it goes from top to bottom. As the complexity of the visual concepts increases, we observe that all methodologies perform poorly in maintaining the subject-fidelity. This applies to $\lambda$-ECLIPSE as well. Interestingly, $\lambda$-ECLIPSE retains the compositional understanding while baselines perform poorly in all scenarios.
In addition, we demonstrate some examples of how P-T2I methods exhibit inconsistency in their generated results depending on each trial. Figure 11 exhibits consistency in single and multiple concept generation across both models. However, Kosmos-G exhibits inconsistency in generating multiple concepts, sometimes placing the Ironman suit in various parts of the dog or omitting it entirely. This also means that $\lambda$-ECLIPSE reduces the diversity of the generated images to improve the consistency. However, this behavior is observed across the UnCLIP family of models.
In our final exploration, we examine the preservation of actual human facial characteristics in target scenarios, combined with varying captions. While each model attempts to retain the original face, none can precisely replicate the exact personal facial features. These examples generally fail to accurately represent the intended compositions, except in one instance in IP-Adapter FaceID.
Appendix F Limitations
Our work represents a pioneering effort in harnessing a method that utilizes the latent space of pre-trained CLIP models for P-T2I generation, it is important to acknowledge certain limitations. Firstly, CLIP, despite its capabilities, does not perfectly capture hierarchical representations, and this limitation can occasionally result in suboptimal outcomes. Consequently, the CLIP contrastive loss can cause the model to deviate from the original subject features, making performing P-T2I on complex concepts, such as human faces, more challenging. We anticipate that the enhancement of CLIP representations will correspondingly improve the performance of our framework in the realm of P-T2I mapping. The $\lambda$-ECLIPSE model has been trained with 34 million parameters and 1.6 million images. However, increasing the data quality and parameters might lead to improvements in the results.
Appendix G Broader Impact
The recent retraction of the LAION-5B dataset, prompted by concerns over the presence of data samples potentially breaching ethical guidelines, has significantly affected various research initiatives. This situation underscores the crucial importance of careful dataset selection and raises a cautionary note for users about the ethical dimensions involved. In line with this, the community needs caution in the application of associated methodologies as well. Subject-driven image generation or Personalized Text-to-Image (P-T2I) methods have the potential to be a transformative tool in numerous domains.
@@ Appendix G Broader Impact
For their positive influence, they enable users to effortlessly generate, modify, and synthesize original subjects into diverse environments, thereby enriching creative expression. On the other hand, the ease of altering and creating images raises concerns about the responsible use of this technology which requires significant ethical and legal considerations. We recommend developers provide a more secure way (such as image attribution) for end-users to ensure accountability for misuse of such models [18]. Users must be acutely aware of being able to infringe intellectual property rights and create misleading or harmful content. As such, those employing subject-driven image-generation techniques should exercise careful judgment, ensuring that their work adheres to ethical standards and legal boundaries. It is imperative that the broader implications of this technology are considered, and that a commitment to responsible and conscientious use guides its application.
@@
Figure 10: Qualitative examples of increasing complexity of novel visual concepts as we move from top to bottom.

### 09 Apr 2024 v2

@@ Acknowledgments
This work was supported by NSF RI grants #1750082, #2132724, and CPS grant #2038666. We thank the Research Computing (RC) at Arizona State University (ASU) for providing computing resources. The views and opinions of the authors expressed herein do not necessarily state or reflect those of the funding agencies and employers.
@@ Appendix A Preliminaries for T2I Diffusion Models
As evidenced in numerous contemporary studies regarding T2I models, Stable Diffusion (SD) [38] has emerged as a predominant backbone for T2I models. SD involves training diffusion models in latent space, reversing a forward diffusion process that introduces noise into the image. A notable feature of SD is its integration of cross-attention, facilitating various conditions like text input. Operating in VQ-VAE latent space, SD not only achieves exceptional generative performance surpassing that in pixel space but also significantly reduces the computational demands.
UnCLIP models (such as DALL-E 2) are very similar to the Stable Diffusion. In contrast, the UnCLIP takes the modular approach. UnCLIP first trains the diffusion text-to-image to the image prior ($f_{\theta}$) to estimate the image embeddings ($z_{x}$) from the text embeddings ($z_{y}$). In parallel, a UNet-like diffusion image generator ($h_{\phi}$) is trained to generate images ($x$) conditioned on ground truth vision embeddings ($z_{x}$).
@@
$\mathcal{L}_{prior}=\underset{\begin{subarray}{c}t\sim[0,T],\\ z_{x}^{(t)}\sim q(t,z_{x})\end{subarray}}{\mathbb{E}}\Big{[}||z_{x}-f_{\theta}(z_{x}^{(t)},t,z_{y})||_{2}^{2}\Big{]}.$ (3)
ECLIPSE proposes the contrastive learning strategy (Eq. 1 – main paper) instead of minimizing Eq. 3. The diffusion image generator is trained by following standard $\epsilon$-prediction formulation. Here, $h_{\phi}$ will estimate the ground truth added Gaussian noise $\epsilon\sim N(0,I)$, given the noise image $X^{(t)}$ for all timesteps $t\sim[0,T]$ and input conditions (such as $z_{x}$, $z_{y}$).
$\mathcal{L}_{decoder}=\underset{\begin{subarray}{c}\epsilon\sim N(0,I)\\ t\sim[0,T],\\ (z_{x},~{}z_{y})\end{subarray}}{\mathbb{E}}\Big{[}||\epsilon-h_{\phi}(x^{(t)},t,z_{x},z_{y})||_{2}^{2}\Big{]}.$ (4)
@@ Appendix B Image-Text Interleaved Training Details
Figure 8: Examples of image-text interleaved training data. The left column shows the input of the prior model and the right images shows the ground truth corresponding images. Note: these examples are generated from $\lambda$-ECLIPSE for better interpretability.
@@
Appendix C Implementation Details
Appendix D $\lambda$-ECLIPSE with Finetuning
Figure 9: DreamBooth (Stable Diffusion v1.5) vs. $\lambda$-ECLIPSE  (with fine-tuning) w.r.t. DINO and CLIP-T metrics on Dreambench.
As demonstrated in the main paper (Table 2), the superiority of fine-tuning-based personalization methodologies, whether applied to single-subject or multi-subject frameworks, over non-fine-tuning alternatives is evident. Consequently, we have expanded our analysis through additional fine-tuning of the $\lambda$-ECLIPSE.
Experimental Setup.
Given that $\lambda$-ECLIPSE effectively trains the T2I prior, capturing concept-specific features to a significant degree, we opted not to further optimize this component. Our focus shifted to exclusively fine-tuning the diffusion UNet model ($h_{\phi}$), employing the AdamW optimizer at a learning rate of 1e-5, without warm-up steps. For the DreamBooth application within the Stable Diffusion v1.5 model, we selected a learning rate of 5e-6, maintaining consistency in other hyperparameters. To simplify, we excluded the use of a prior preservation regularizer and conducted training on the Dreambench platform using a single RTX A6000 GPU.
Results.
Our findings, illustrated in Figure 9, reveal that $\lambda$-ECLIPSE and DreamBooth exhibit improved performance with incremental fine-tuning steps. Notably, the DINO score improved from 0.61 to 0.68 with few optimization steps and outperforms the baselines (see Table 2). A detailed analysis indicates that while DreamBooth’s DINO score improves, its CLIP-T performance diminishes, hinting at concept overfitting. Conversely, $\lambda$-ECLIPSE consistently improves in DINO scoring without adversely impacting the CLIP-T performance, underscoring the efficacy of our image-text interleaved training approach at the prior stage. Qualitative comparisons, as shown in Figure 10, further highlight the benefits of fine-tuning $\lambda$-ECLIPSE with minimal steps.
Advantages of fine-tuning $\lambda$-ECLIPSE.
The fine-tuning of $\lambda$-ECLIPSE, in comparison to the baselines, reveals two key benefits: 1) Achieving state-of-the-art (SOTA) performance within a few finetuning steps. 2) Unlike the Stable Diffusion model, which exhibits catastrophic forgetting of nearby concepts post-DreamBooth fine-tuning, $\lambda$-ECLIPSE maintains previous knowledge. This suggests that a single model is sufficient to effectively fine-tune across multiple concepts together.
This analysis underscores the strategic advantages and enhanced efficiency of fine-tuning $\lambda$-ECLIPSE for personalized applications in complex visual data processing.
Appendix E Extended P-T2I Baselines Comparison
@@
Method Multi-Subject Finetuning-Free Base-Model # of Input Images
@@
We further expand our comparative analysis of P-T2I methods encompassing a total of 33 approaches including ours and parallel works. Table 7 summarizes them into four crucial aspects: 1) multi-subject support, 2) fine-tuning free, 3) base model types, and 4) the required number of input images. To summarize, $\lambda$-ECLIPSE is the only methodology built on top of the UnCLIP models while supporting multi-subject driven image generation with fine-tuning free, and only requires a single reference image input for the training. We detail the comparison below:
Multi-Subject Generation.
Multi-subject generation enables users to integrate multiple personal subjects to generate an image that follows the text prompts and aligns with all the concept visuals. In total, 15 of the 33 methods offer this capability, while 6 methods support fast multi-subject personalization, others demand separate training for each subject to be learned and then an additional fusing step for combining the learned subjects is required (i.e. Zip-LoRA, Mix-of-Show). Among these methods, only a few can learn auxiliary guided information such as canny edge, depth maps, or open-pose and adapt style variation (i.e. Kosmos-G).
@@
Many methods require test-time fine-tuning. Each varies on which part alteration occurs, as early models tend to modify the whole UNet. In contrast, recent models tune a small portion of the cross-attention layers or introduce additional layers performing as adapters. In our analysis of P-T2I methodologies, 14 out of 33 methods employ a finetuning-free approach which enables fast personalization.
@@
A majority of the reviewed models utilize diffusion models, with Stable Diffusion being the predominant choice, spanning versions 1.4, 1.5, 2.1, and XL. Few adapt Imagen (SuTi, Taming) and Mix-of-show employs ChillOutMix as their pre-trained model, known for its adeptness at preserving realistic concepts like human faces. A unique outlier in this landscape is our $\lambda$-ECLIPSE, the only one that eschews the use of any diffusion prior model.
Easiness of Use.
@@
A more user-friendly model typically requires a single reference image per subject, as opposed to multiple images of the same subject. In our study, 19 methods offer P-T2I capabilities with just one input image. In contrast, others often require 4 to 5 images of the subject. Additionally, some methods necessitate storage space for learned concepts, ranging from a few hundred kilobytes (e.g., Perfusion, HyperDreamBooth) to several megabytes (e.g., Zip-LoRA). Our method stands out by eliminating the need for individual concept pre-learning or storing any artifacts for P-T2I utilization, offering a streamlined, efficient user experience.
Appendix F Multibench Dataset
Table 8: Example of prompt templates used for Multibench dataset. Subjects presented in Table 9 are placed in {}.
Two subjects Three subjects
{} in the {} {} with a {} and {}
{} wearing a {} {} is playing with {} in {}
{} chasing a {} {} with {} in front of {}
{} looking at a {} {} with a {} and a view of the {}
{} is sitting on a {} {} with a {} and {} in the background
{} standing on a {}
{} and {} playing in the garden
{} and {} on top of the mountain
{} and {} in the jungle
{} and {} in the snow
{} and {} on the beach
{} and {} on a cobblestone street
{} and {} standing next to each other
Table 9: Number of occurrences of unique subject categories. The left side of the table are subjects used for two subjects prompts, and the right side of the table are subjects used for three subjects prompts.
Two subjects Three subjects
dog 76 boat 5 dog 81 rainbow 35
cat 76 park 4 stuffed animal 105 ruins 35
bird 76 ruins 9 toy 105 tower 35
horse 73 castle 5 cat 81 horse 81
guinea pig 73 desert 4 desert 60 bird 81
glasses 5 rainbow 5 hill 60 guinea pig 81
hat 5 candle 5 castle 45 guitar 25
tower 10 backpack 3 backpack 65 french horn 25
can 130 vase 25
candle 65 robot 25
church 35
We provide additional qualitative results in Figure 5. For the multi-subject image benchmark, our dataset comprises 2,308 unique prompts, segmented into 904 two-subject and 1,476 three-subject prompts. This dataset integrates subjects from the original DreamBench dataset, featuring 30 distinct concepts. We expanded the dataset by incorporating additional concepts vital for two and three subject-specific prompts, such as various parks, hats, glasses, and more. Prompt templates and the count of unique subject categories featured in prompts are detailed in Tables 8 and  9, respectively. Overall, the dataset includes 217 two-subject compositions and 405 three-subject compositions, enriching the benchmark’s diversity and comprehensiveness.
Appendix G Qualitative Results & Failure Cases
In this section, we showcase a collection of detailed qualitative examples from the P-T2I generation process, highlighting the challenges of crafting complex compositions within $\lambda$-ECLIPSE and comparative models. As depicted in Figure 11, the complexity of the showcased examples progressively increases, illustrating a noticeable escalation in the intricacy of visual concepts from the top to the bottom of the figure. With the rising complexity, we note a universal decline in the ability of all methodologies, including $\lambda$-ECLIPSE, to preserve subject fidelity accurately. Interestingly, despite these challenges, $\lambda$-ECLIPSE demonstrates a better grasp of compositional integrity, unlike the baseline models which falter across all complexity levels.
Moreover, we present instances demonstrating the variability in outcomes produced by P-T2I methods across different trials. As illustrated in Figure 12, while there is a semblance of consistency in generating single and multiple concepts between models, Kosmos-G specifically shows variability in rendering multiple concepts—occasionally misplacing elements of the Ironman suit on a dog or failing to include it altogether. This phenomenon suggests that $\lambda$-ECLIPSE minimizes image diversity to enhance result consistency, a trait observed across the UnCLIP model family.
Figure 10 offers qualitative insights into the performance of $\lambda$-ECLIPSE without and with minimal fine-tuning. It is evident that in certain edge cases, where $\lambda$-ECLIPSE initially struggles to fully grasp novel visual concepts without finetuning, a modest application of few optimization iterations significantly enhances concept capture. Further optimization not only preserves text composition but also enriches minor, subject-specific details, underscoring the adaptability and finesse of $\lambda$-ECLIPSE in nuanced image generation.
Moreover, in our evaluations using the Multibench dataset, we noticed that both the baseline models (Kosmos-G and Emu2) and $\lambda$-ECLIPSE encounter difficulties in precisely maintaining all subject-specific details, as depicted in Figure 14. This underscores that zero-shot multi-subject P-T2I generation remains a significant challenge in the field. Further, we explored how well each model preserves genuine human facial characteristics in various scenarios, particularly when combined with differing captions. The qualitative examples in Figure 13 shed light on this aspect. Although each model strives to maintain the original facial features, none succeeds in replicating the specific personal facial details accurately. These instances typically fall short of precisely conveying the intended compositions, with the exception of one scenario in IP-Adapter FaceID, indicating a notable area for future improvements in model performance.
Appendix H Limitations
Our work marks a pioneering venture into leveraging the latent space of pre-trained CLIP models for P-T2I generation. Nonetheless, it’s crucial to recognize certain constraints. Primarily, despite its strengths, CLIP’s inability to perfectly capture hierarchical representations occasionally leads to less-than-ideal results. This issue, stemming from the CLIP contrastive loss, can cause deviations from the original subject features, particularly when generating P-T2I for complex subjects like human faces. We believe that enhancing CLIP’s representations could significantly boost our framework’s efficacy in P-T2I mapping. The $\lambda$-ECLIPSE model, trained on 34 million parameters and 1.6 million images, presents a substantial foundation. Yet, there’s potential for further refinement, as increasing the quality of data and the number of parameters could yield even better outcomes.
Appendix I Broader Impact
@@
Subject-driven image generation or Personalized Text-to-Image (P-T2I) methods have the potential to be a transformative tool in numerous domains. For their positive influence, they enable users to effortlessly generate, modify, and synthesize original subjects into diverse environments, thereby enriching creative expression. On the other hand, the ease of altering and creating images raises concerns about the responsible use of this technology which requires significant ethical and legal considerations. Users must be acutely aware of being able to infringe intellectual property rights and create misleading or harmful content. We recommend developers provide a more secure way such as image attribution [18] for end-users to ensure accountability for misuse of such models. As such, those employing subject-driven image-generation techniques should exercise careful judgment, ensuring that their work adheres to ethical standards and legal boundaries. It is imperative that the broader implications of this technology are considered, and that a commitment to responsible and conscientious use guides its application.
Figure 10: Qualitative examples of $\lambda$-ECLIPSE without finetuning and different stages of finetuning.
@@
Figure 11: Qualitative examples of the increasing complexity of novel visual concepts as we move from top to bottom.
@@
Figure 14: Qualitative examples of showcasing the failure cases on Multibench of Kosmos-G, Emu2, and $\lambda$-ECLIPSE.